# Predictable Scale: Part I — Optimal Hyperparameter Scaling Law in Large Language Model Pretraining

### v4

@@ A.2 Model Structural Parameters And Results
0 2.15E+08 1.14E+10 960 9368 15 7
1 4.29E+08 5.00E+10 1280 9472 10 10
2 2.68E+08 8.00E+10 1024 9552 16 8
3 4.29E+08 8.00E+09 1280 9472 10 10
4 1.07E+09 2.00E+10 2048 8192 16 16
5 5.37E+08 1.00E+10 1280 9048 10 13
6 2.15E+08 4.00E+09 960 9368 15 7
7 2.68E+08 5.00E+09 1024 9552 16 8
8 2.68E+08 1.42E+10 1024 9552 16 8
9 1.07E+09 5.69E+10 2048 8192 16 16
10 2.15E+08 1.00E+11 960 9368 15 7
11 4.29E+08 2.27E+10 1280 9472 10 10
12 5.37E+08 2.84E+10 1280 9048 10 13
13 2.15E+08 2.00E+10 960 9368 15 7
14 4.29E+08 4.00E+10 1280 9472 10 10
15 2.68E+08 2.50E+10 1024 9552 16 8
16 5.37E+08 5.00E+10 1280 9048 10 13
17 1.07E+09 1.00E+11 2048 8192 16 16
@@
0 2150612992 2000000000 1408 11 16 89 352 1 187973632
1 2150612992 2000000000 1408 11 16 88 352 2 232579072
2 2155174912 2000000000 1408 11 16 8 3528 1 590436352
3 2156188672 2000000000 1408 11 16 8 2888 3 1241270272
4 2150612992 4000000000 1408 11 16 89 352 1 187973632
5 2150612992 4000000000 1408 11 16 88 352 2 232579072
6 2155174912 4000000000 1408 11 16 8 3528 1 590436352
7 2156188672 4000000000 1408 11 16 8 2888 3 1241270272
8 2150612992 8000000000 1408 11 16 89 352 1 187973632
9 2150612992 8000000000 1408 11 16 88 352 2 232579072
10 2155174912 8000000000 1408 11 16 8 3528 1 590436352
11 2156188672 8000000000 1408 11 16 8 2888 3 1241270272
12 2150612992 20000000000 1408 11 16 89 352 1 187973632
13 2150612992 20000000000 1408 11 16 88 352 2 232579072
14 2155174912 20000000000 1408 11 16 8 3528 1 590436352
15 2156188672 20000000000 1408 11 16 8 2888 3 1241270272
@@
Figure 11: Illustration of Hyperparameter Configuration Space for Model 0.
Figure 12: Illustration of Hyperparameter Configuration Space for Model 1.
Figure 13: Illustration of Hyperparameter Configuration Space for Model 2.
Figure 14: Illustration of Hyperparameter Configuration Space for Model 3.
Figure 15: Illustration of Hyperparameter Configuration Space for Model 4.
Figure 16: Illustration of Hyperparameter Configuration Space for Model 5.
Figure 17: Illustration of Hyperparameter Configuration Space for Model 6.
Figure 18: Illustration of Hyperparameter Configuration Space for Model 7.
Figure 19: Illustration of Hyperparameter Configuration Space for Model 8.
Figure 20: Illustration of Hyperparameter Configuration Space for Model 9.
Figure 21: Illustration of Hyperparameter Configuration Space for Model 10.
Figure 22: Illustration of Hyperparameter Configuration Space for Model 11.
Figure 23: Illustration of Hyperparameter Configuration Space for Model 12.
Figure 24: Illustration of Hyperparameter Configuration Space for Model 13.
Figure 25: Illustration of Hyperparameter Configuration Space for Model 14.
Figure 26: Illustration of Hyperparameter Configuration Space for Model 15.
Figure 27: Illustration of Hyperparameter Configuration Space for Model 16.
Figure 28: Illustration of Hyperparameter Configuration Space for Model 17.
Figure 29: Illustration of Hyperparameter Configuration Space for MoE Model 0.
Figure 30: Illustration of Hyperparameter Configuration Space for MoE Model 1.
Figure 31: Illustration of Hyperparameter Configuration Space for MoE Model 2.
Figure 32: Illustration of Hyperparameter Configuration Space for MoE Model 3.
Figure 33: Illustration of Hyperparameter Configuration Space for MoE Model 4.
Figure 34: Illustration of Hyperparameter Configuration Space for MoE Model 5.
Figure 35: Illustration of Hyperparameter Configuration Space for MoE Model 6.
Figure 36: Illustration of Hyperparameter Configuration Space for MoE Model 7.
Figure 37: Illustration of Hyperparameter Configuration Space for MoE Model 8.
Figure 38: Illustration of Hyperparameter Configuration Space for MoE Model 9.
Figure 39: Illustration of Hyperparameter Configuration Space for MoE Model 10.
Figure 40: Illustration of Hyperparameter Configuration Space for MoE Model 11.
Figure 41: Illustration of Hyperparameter Configuration Space for MoE Model 12.
Figure 42: Illustration of Hyperparameter Configuration Space for MoE Model 13.
Figure 43: Illustration of Hyperparameter Configuration Space for MoE Model 14.
Figure 44: Illustration of Hyperparameter Configuration Space for MoE Model 15.

### v6

@@ A.3 Model Structural Parameters
1 $2.15\times 10^{8}$ $1.14\times 10^{10}$ 960 9368 15 7
2 $4.29\times 10^{8}$ $5.00\times 10^{10}$ 1280 9472 10 10
3 $2.68\times 10^{8}$ $8.00\times 10^{10}$ 1024 9552 16 8
4 $4.29\times 10^{8}$ $8.00\times 10^{9}$ 1280 9472 10 10
5 $1.07\times 10^{9}$ $2.00\times 10^{10}$ 2048 8192 16 16
6 $5.37\times 10^{8}$ $1.00\times 10^{10}$ 1280 9048 10 13
7 $2.15\times 10^{8}$ $4.00\times 10^{9}$ 960 9368 15 7
8 $2.68\times 10^{8}$ $5.00\times 10^{9}$ 1024 9552 16 8
9 $2.68\times 10^{8}$ $1.42\times 10^{10}$ 1024 9552 16 8
10 $1.07\times 10^{9}$ $5.69\times 10^{10}$ 2048 8192 16 16
11 $2.15\times 10^{8}$ $1.00\times 10^{11}$ 960 9368 15 7
12 $4.29\times 10^{8}$ $2.27\times 10^{10}$ 1280 9472 10 10
13 $5.37\times 10^{8}$ $2.84\times 10^{10}$ 1280 9048 10 13
14 $2.15\times 10^{8}$ $2.00\times 10^{10}$ 960 9368 15 7
15 $4.29\times 10^{8}$ $4.00\times 10^{10}$ 1280 9472 10 10
16 $2.68\times 10^{8}$ $2.50\times 10^{10}$ 1024 9552 16 8
17 $5.37\times 10^{8}$ $5.00\times 10^{10}$ 1280 9048 10 13
18 $1.07\times 10^{9}$ $1.00\times 10^{11}$ 2048 8192 16 16
@@
1 $2.151\times 10^{9}$ $2.00\times 10^{9}$ 1408 11 16 89 352 1 $1.88\times 10^{8}$
2 $2.151\times 10^{9}$ $2.00\times 10^{9}$ 1408 11 16 88 352 2 $2.33\times 10^{8}$
3 $2.155\times 10^{9}$ $2.00\times 10^{9}$ 1408 11 16 8 3528 1 $5.90\times 10^{8}$
4 $2.156\times 10^{9}$ $2.00\times 10^{9}$ 1408 11 16 8 2888 3 $1.24\times 10^{9}$
5 $2.151\times 10^{9}$ $4.00\times 10^{9}$ 1408 11 16 89 352 1 $1.88\times 10^{8}$
6 $2.151\times 10^{9}$ $4.00\times 10^{9}$ 1408 11 16 88 352 2 $2.33\times 10^{8}$
7 $2.155\times 10^{9}$ $4.00\times 10^{9}$ 1408 11 16 8 3528 1 $5.90\times 10^{8}$
8 $2.156\times 10^{9}$ $4.00\times 10^{9}$ 1408 11 16 8 2888 3 $1.24\times 10^{9}$
9 $2.151\times 10^{9}$ $8.00\times 10^{9}$ 1408 11 16 89 352 1 $1.88\times 10^{8}$
10 $2.151\times 10^{9}$ $8.00\times 10^{9}$ 1408 11 16 88 352 2 $2.33\times 10^{8}$
11 $2.155\times 10^{9}$ $8.00\times 10^{9}$ 1408 11 16 8 3528 1 $5.90\times 10^{8}$
12 $2.156\times 10^{9}$ $8.00\times 10^{9}$ 1408 11 16 8 2888 3 $1.24\times 10^{9}$
13 $2.151\times 10^{9}$ $2.00\times 10^{10}$ 1408 11 16 89 352 1 $1.88\times 10^{8}$
14 $2.151\times 10^{9}$ $2.00\times 10^{10}$ 1408 11 16 88 352 2 $2.33\times 10^{8}$
15 $2.155\times 10^{9}$ $2.00\times 10^{10}$ 1408 11 16 8 3528 1 $5.90\times 10^{8}$
16 $2.156\times 10^{9}$ $2.00\times 10^{10}$ 1408 11 16 8 2888 3 $1.24\times 10^{9}$
17 $6.510\times 10^{9}$ $1.00\times 10^{10}$ 2048 16 24 82 512 2 $7.26\times 10^{8}$
18 $6.510\times 10^{9}$ $1.30\times 10^{11}$ 2048 16 24 82 512 2 $7.26\times 10^{8}$
@@
A.4 Composition of Training Datasets
Tab. 7 details the dataset weight percentages for four training recipes: Baseline, Code+Math, More Code+Math, and En-CN (English-Chinese bilingual).
@@ A.4 Composition of Training Datasets
A.5 Statistical Validation of Batch Size Scaling Relationships
Empirical validation.
This part provides a statistical analysis to examine the claim in the main text that batch size $B$ is independent of model parameter count $N$ but dependent on training dataset size $D$. We conducted a multivariate regression analysis on all experimental configurations. All variables are log-transformed to linearize power-law relationships (logB, logN, logD), consistent with the main text. Three regression formulations were compared:
N-only Formulation: $\log B=\beta_{0}+\beta_{1}\log N$
D-only Formulation: $\log B=\beta_{0}+\beta_{2}\log D$
Full Formulation: $\log B=\beta_{0}+\beta_{1}\log N+\beta_{2}\log D$
We fit each of these models using ordinary least squares (OLS) regression and perform hierarchical F-tests to assess the contribution of $\log N$ and $\log D$ to predicting $\log B$. The results are summarized in Table  8 and Table 9 .
Table 8: Regression formulation comparisons
Comparison Adj. $R^{2}$ $\Delta R^{2}$ VS Full F-statistic
N-only -0.032 -85.4% 0.08
D-only 0.821 -0.2% 138.2
Full 0.823 0.0% 70.59
Table 9: Coefficient analysis for Full formulation
Predictor Coeff. Std. Error t-value p-value 95% CI
$\log N$0.087 0.075 -1.158 0.257 [-0.241, 0.067]
$\log D$ 0.580 0.049 11.863 $<$0.001 [0.480, 0.680]
As shown in Appendix Table 8 and Table 9, the D-only formulation achieves nearly identical explanatory power as the full model ($R^{2}=0.821$ vs. $0.823$), while the N-only model performs poorly ($R^{2}<0$). Moreover, in the full model, the coefficient of $\log D$ is highly significant ($p<0.001$), whereas that of $\log N$ is not ($p=0.257$). These results confirm that the optimal batch size scales with $D$, but not with $N$.
A.6 Loss Landscape Convexity
Experiments are carried out on 18 unique model architectures, with their precise parameter settings summarized in Appendix 5. For each model, the Smooth Loss metric is derived through a grid search over the hyperparameters of learning rate and batch size. Visualization of the smooth loss landscapes substantiates the claims in the main text, demonstrating a consistent convex relationship between the hyperparameters (learning rate and batch size) and the training loss across all models. These findings confirm the robustness and stability of the optimization process as asserted in the primary discussion.
(a) Model 1
(b) Model 2
Figure 10: Illustration of Hyperparameter Configuration Spaces for Models 1 and 2.
(a) Model 3
(b) Model 4
Figure 11: Illustration of Hyperparameter Configuration Spaces for Models 3 and 4.
(a) Model 5
(b) Model 6
Figure 12: Illustration of Hyperparameter Configuration Spaces for Models 5 and 6.
(a) Model 7
(b) Model 8
Figure 13: Illustration of Hyperparameter Configuration Spaces for Models 7 and 8.
(a) Model 9
(b) Model 10
Figure 14: Illustration of Hyperparameter Configuration Spaces for Models 9 and 10.
(a) Model 11
(b) Model 12
Figure 15: Illustration of Hyperparameter Configuration Spaces for Models 11 and 12.
(a) Model 13
(b) Model 14
Figure 16: Illustration of Hyperparameter Configuration Spaces for Models 13 and 14.
(a) Model 15
(b) Model 16
Figure 17: Illustration of Hyperparameter Configuration Spaces for Models 15 and 16.
(a) Model 17
(b) Model 18
Figure 18: Illustration of Hyperparameter Configuration Spaces for Models 17 and 18.
A.7 Dense Models
This section presents the full set of hyperparameter configuration space visualizations for the 18 dense Transformer models described in Table 5. Each plot illustrates the validation loss surface as a function of learning rate (LR) and batch size (BS), using a log–log scale for both axes. These visualizations reveal consistent trends across model scales, including the emergence of convex, bowl-shaped minima and smooth shifts in optimal hyperparameter regions. They serve as empirical evidence for the predictive structure captured by the scaling law, and demonstrate its robustness across a wide range of dense model sizes.
(a) Model 1
(b) Model 2
Figure 19: Illustration of Hyperparameter Configuration Spaces for Models 1 and 2.
(a) Model 3
(b) Model 4
Figure 20: Illustration of Hyperparameter Configuration Spaces for Models 3 and 4.
(a) Model 5
(b) Model 6
Figure 21: Illustration of Hyperparameter Configuration Spaces for Models 5 and 6.
(a) Model 7
(b) Model 8
Figure 22: Illustration of Hyperparameter Configuration Spaces for Models 7 and 8.
(a) Model 9
(b) Model 10
Figure 23: Illustration of Hyperparameter Configuration Spaces for Models 9 and 10.
(a) Model 11
(b) Model 12
Figure 24: Illustration of Hyperparameter Configuration Spaces for Models 11 and 12.
(a) Model 13
(b) Model 14
Figure 25: Illustration of Hyperparameter Configuration Spaces for Models 13 and 14.
(a) Model 15
(b) Model 16
Figure 26: Illustration of Hyperparameter Configuration Spaces for Models 15 and 16.
(a) Model 17
(b) Model 18
Figure 27: Illustration of Hyperparameter Configuration Spaces for Models 17 and 18.
A.8 MoE Models
To assess the generality of HP scaling laws beyond dense Transformers, we conduct a comprehensive study on MoE models, which activate only a subset of experts per token. We evaluate 16 distinct configurations (see Table 6), varying both total parameter count and sparsity. For each model, we sweep learning rate (LR) and batch size (BS) over the same logarithmic grid used in our dense experiments.
(a) MoE Model 1
(b) MoE Model 2
Figure 28: Illustration of Hyperparameter Configuration Spaces for MoE Models 1 and 2.
(a) MoE Model 3
(b) MoE Model 4
Figure 29: Illustration of Hyperparameter Configuration Spaces for MoE Models 3 and 4.
(a) MoE Model 5
(b) MoE Model 6
Figure 30: Illustration of Hyperparameter Configuration Spaces for MoE Models 5 and 6.
(a) MoE Model 7
(b) MoE Model 8
Figure 31: Illustration of Hyperparameter Configuration Spaces for MoE Models 7 and 8.
(a) MoE Model 9
(b) MoE Model 10
Figure 32: Illustration of Hyperparameter Configuration Spaces for MoE Models 9 and 10.
(a) MoE Model 11
(b) MoE Model 12
Figure 33: Illustration of Hyperparameter Configuration Spaces for MoE Models 11 and 12.
(a) MoE Model 13
(b) MoE Model 14
Figure 34: Illustration of Hyperparameter Configuration Spaces for MoE Models 13 and 14.
(a) MoE Model 15
(b) MoE Model 16
Figure 35: Illustration of Hyperparameter Configuration Spaces for MoE Models 15 and 16.
A.9 Limitations